Two cars, A and B, are traveling in the same direction, although car A is 187 m behind car B. The speed of A is 22.4 m/s, and the speed of B is 17.8 m/s. How much time does it take for A to catch B?
Answer #1

Relative velocity, v = Va - Vb = 22.4 - 17.8 = 4.6 m/s

Relative position, x = 187 m

Time, t = x/v = 187/4.6 = 40.65 sec

Comment in case any doubt please rate my answer......